# MATHGLANCE: Multimodal Large Language Models Do Not Know Where to Look in Mathematical Diagrams

### v1

@@ Abstract
Diagrams serve as a fundamental form of visual language, representing complex concepts and their interrelationships through structured symbols, shapes, and spatial arrangements. Unlike natural images, their inherently symbolic and abstract nature poses significant challenges for Multimodal Large Language Models (MLLMs). However, current benchmarks conflate perceptual and reasoning tasks, making it difficult to assess whether MLLMs genuinely understand mathematical diagrams beyond superficial pattern recognition. To address this gap, we introduce MATHGLANCE, a benchmark specifically designed to isolate and evaluate mathematical perception in MLLMs. MATHGLANCE comprises 1.2K images and 1.6K carefully curated questions spanning four perception tasks: shape classification, object counting, relationship identification, and object grounding, covering diverse domains including plane geometry, solid geometry, and graphical representations. Our evaluation of MLLMs reveals that their ability to understand diagrams is notably limited, particularly in fine-grained grounding tasks. In response, we construct GeoPeP, a perception-oriented dataset of 200K structured geometry image-text pairs explicitly annotated with geometric primitives and precise spatial relationships. Training MLLM on GeoPeP leads to significant gains in perceptual accuracy, which in turn substantially improves mathematical reasoning. Our benchmark and dataset establish critical standards for evaluating and advancing multimodal mathematical understanding, providing valuable resources and insights to foster future MLLM research.
Figure 1: Performance comparison of current Multimodal Large Language Models (MLLMs) on MATHGLANCE reveals a stark contrast: while humans can solve fundamental perception tasks “at a glance”, these tasks remain highly challenging for MLLMs—particularly in fine-grained grounding. SVE-Math-DeepSeek $\!\!{}^{+}$-7B, trained with the proposed perception-oriented GeoPeP, shows significantly enhanced perceptual capabilities.
@@ 1 Introduction
Computer vision research has traditionally focused on developing models that can perform fundamental perception tasks [17, 30] such as object detection [36, 56], segmentation [28, 37], and spatial reasoning [12]. These capabilities form the foundation for higher-level visual reasoning and decision-making. However, recent advances in Multimodal Large Language Models (MLLMs) [1, 32, 42] have shifted the paradigm toward vision-language integration, where visual information is processed through natural language interfaces rather than specialized perception modules.
Figure 2: Illustration of diagram-caption-based alignment training datasets, w.r.t., AutoGeo, MAVIS, and our proposed GeoPeP (Fig. 2(a)). Fig. 2(b) demonstrates a positive correlation between low-level perception and high-level reasoning tasks, evaluated on MATHGLANCE and MathVista. Clear diagram perception leads to substantial improvements in mathematical reasoning performance.
Despite their remarkable successes in general vision tasks, current MLLMs face considerable challenges in accurately interpreting mathematical diagrams—abstract visual representations characterized by precise geometric structures and symbolic notations [34, 53]. These diagrams provide a unique perceptual challenge: the explicit structural constraints require genuine visual interpretation rather than pattern recognition or textual heuristics, positioning them as an essential benchmark for evaluating the true visual perception capabilities of multimodal models. Such diagrams are pervasive across educational contexts, scientific communication, and problem-solving in STEM disciplines. Developing models capable of understanding symbolic information in diagrams, is a critical milestone in advancing machine intelligence  [14, 13].
@@
Recent benchmarks like MathVista [34] and MathVerse [53] have attempted to evaluate mathematical visual reasoning in MLLMs. However, these works conflate perception with reasoning by assessment on complex reasoning tasks that combine diagram interpretation with numerical computation and proof generation. Thus, it remains unclear whether the performance truly reflects the models’ ability to comprehensively understand the symbolic information in diagrams. Misinterpretations at the perception level propagate errors downstream, resulting in incorrect reasoning outcomes and frequent model hallucinations [7, 44, 23].
@@
To address this critical gap, we design a benchmark to isolate and rigorously evaluate mathematical perception in MLLMs. Our MATHGLANCE features problems that humans can solve “at a glance” without extensive reasoning. The benchmark contains 1,198 images and 1,609 carefully designed perception-oriented questions across four distinct task categories: shape classification, object counting, relationship identification, and object grounding. These tasks span three core mathematical domains—plane geometry, solid geometry, and graphical representations (line, bar, and pie graphs)—and question formats (multiple-choice, true/false, and free-form).
Figure 3: Sampled MATHGLANCE  examples from plane geometry w.r.t. each question-answer (Q&A) type, covering shape classification, object counting, relationship identification, and object grounding (from left to right). The green dotted bounding boxe is shown for illustration purposes only and are not provided as input to the models.
Through comprehensive evaluation of current MLLMs on MATHGLANCE, we explore the following questions:
  1. 1.
@@
    Do current MLLMs genuinely perceive mathematical diagrams? Fig. 1 presents the performance of eight current MLLMs—comprising six generic models (both open-source and proprietary) and two math-specific model—across four perception-focused tasks in MATHGLANCE (see Tab. 2 for more results). Among the generic MLLMs, Qwen2.5-VL-7B [6] achieves the highest average accuracy, followed by GPT-4o. But, those models exhibit severe hallucination issues, often responding to simple questions with unnecessarily long chain-of-thought (CoT) reasoning and generating irrelevant image content (see Fig.6). The math-specific SVE-Math-DeepSeek [55], which incorporates a domain-specific visual encoder trained on multi-task geometric primitive detection, achieves comparable performance to Qwen2.5-VL-7B on shape classification and relationship identification tasks. Nevertheless, all evaluated models struggle significantly with fine-grained bounding box grounding. Even those trained on large-scale grounding annotations with nearly 1T natural images—such as Qwen2.5-VL-7B and DeepSeek-VL2-Small—achieve less than 20% accuracy on this task.
    To alleviate this issue, we propose GeoPeP, a Geomatic PercePtion-oriented dataset. Unlike existing image-caption-based visual training datasets (e.g., MAVIS [54], AutoGeo [21]), GeoPeP provides explicit structured diagram information, including shapes, attributes, locations, and relationships (see Fig.2(a)). We first train only the vision-language projector of SVE-Math-DeepSeek using this alignment dataset. Then, we construct instruction conversational data for a second self-fine-tuning (SFT) stage, jointly updating both the projector and the large language model. The resulting model, SVE-Math-DeepSeek $\!{}^{+}$, achieves the highest performance among both generic and math-specific models, particularly excelling in grounding tasks with a +79% improvement.
  2. 2.
    Does stronger perceptual ability lead to better reasoning performance? We hypothesize that the more a model can “see”, the more it can “reason” —that is, clear diagram perception would support stronger reasoning ability. As shown in Fig.2(b), training with our proposed alignment dataset significantly improves perception performance compared to AutoGeo and MAVIS (by 23.0% and 35.6%, respectively). In contrast, the other two variants perform worse than the baseline (SVE-Math-DeepSeek-7B), particularly with MAVIS diagram-caption pairs, due to high uncertainty and out-of-distribution nature of real-world geometric diagrams.

### v2

# Math Blind: Failures in Diagram Understanding Undermine Reasoning in MLLMs

@@ Abstract
Diagrams represent a form of visual language that encodes abstract concepts and relationships through structured symbols and their spatial arrangements. Unlike natural images, they are inherently symbolic, and entirely artificial. They thus pose unique challenges for Multimodal Large Language Models (MLLMs) distinct from natural image processing. Recent studies have shown that MLLMs often exhibit flawed reasoning and hallucinations when handling diagram inputs. We investigate here whether these limitations stem from shortcomings in the models’ ability to interpret diagrams themselves. To this end, we develop a diagnostic test suite that isolates perception from reasoning. Our systematic evaluation reveals that MLLMs perform poorly on basic perceptual tasks, e.g., shape classification, object counting, relationship identification, and object grounding, with near-zero accuracy on fine-grained grounding. Further analysis shows that weak diagram perception leads to “blind faith in text", where models rely on textual shortcuts rather than visual understanding (that is, they are Math Blind). We hypothesize that enabling models to capture the inherent structural properties of diagrams, represented as graphs of primitives and their interrelationships, is essential for improving diagram understanding. Experiments with 7B and 32B MLLMs validate this assumption, with models trained on such representations achieving a +79% gain on the grounding task. Crucially, these gains transfer to reasoning, achieving 3–4% cross-suite improvements on three public benchmarks even without additional chain-of-thought reasoning data. Our findings demonstrate that low-level perception supports faithful high-level reasoning in mathematical MLLMs. We provide both methodological frameworks and empirical evidence to guide future research in this direction.
[…]
@@ 1 Introduction
Computer vision has traditionally focused on image-based perception tasks such as object detection [1, 2], segmentation [3, 4], and spatial reasoning [5, 6, 7]. These capabilities form the foundation for higher-level visual reasoning in recent advances of Multimodal Large Language Models (MLLMs) [8, 9, 10]. Despite remarkable successes in general vision tasks, current MLLMs face considerable challenges in interpreting mathematical diagrams.
Although both natural images and symbolic diagrams can be represented as grids of pixels, they constitute very different forms of information. Images represent samples of the intensity of the real world. Diagrams, despite taking many forms, are uniformly generated by humans as abstract visual representations of concepts characterized by precise geometric structures and symbolic notations [11, 12]. Images are natural collections of signals, where diagrams are artificial collections of symbols. Diagrams are critical to human visual communication across educational contexts, scientific discourse, and STEM problem-solving, but the problem of analyzing symbolic visual information is far broader.
@@
Recent benchmarks such as MathVista [11] and MathVerse [12] evaluate mathematical visual reasoning in MLLMs. These works conflate perception with reasoning, however, because they assess performance on complex tasks that combine diagram interpretation and mathematical reasoning. It thus remains unclear whether the performance truly reflects the models’ ability to comprehensively understand the symbolic information in diagrams. Perceptual misinterpretations propagate downstream, leading to faulty reasoning and frequent hallucinations, although the final answer may occasionally still be correct [13, 14, 15]. Developing models capable of understanding symbolic information in diagrams is a critical milestone in advancing machine intelligence [16, 17].
Figure 1: Performance on MATHEMETRIC reveals that diagram interpretation is challenging for MLLMs, particularly in fine-grained grounding tasks that require precise spatial localization. SVE-Math-DeepSeek+-7B, trained on GEOMETRIC, significantly outperforms comparators, validating the structure-aware geometric data design.
@@
Motivated by the above, we have designed a diagnostic benchmark to isolate and rigorously evaluate mathematical perception in MLLMs. MATHEMETRIC features problems that humans can solve “at a glance” without extensive reasoning. The benchmark contains 1,198 images and 1,609 carefully designed perception-oriented questions across four distinct task categories: shape classification, object counting, relationship identification, and object grounding. These tasks span three core mathematical domains—plane geometry, solid geometry, and graphical representations (line, bar, and pie graphs)—and question formats (multiple-choice, true/false, and free-form).
We systematically evaluate current MLLMs on MATHEMETRIC to investigate key questions:
@@
1. Do current MLLMs genuinely perceive mathematical diagrams? Fig. 1 presents the performance of eight current MLLMs—six generic models and two math-specific models—across four perception-focused tasks in MATHEMETRIC (see Tab. 1 for more MLLM results). Among the generic MLLMs, Qwen2.5-VL-7B [18] achieves the highest average accuracy, followed by GPT-4o. But both suffer from severe hallucinations—responding to simple questions with unnecessarily long chains-of-thought and generating irrelevant visual content (see Fig. 6). The math-specific SVE-Math-DeepSeek [19], with a geometric primitive visual encoder, performs on par with Qwen2.5-VL-7B in shape classification and relationship identification. However, all models, including Qwen2.5-VL-7B and DeepSeek-VL2-Small, which are trained on nearly 2T natural images, consistently underperform on fine-grained bounding box grounding, with accuracy below 20%.
We then quantitatively analyze factors influencing MLLMs’ perceptual capabilities in §3.3, providing more convincing evidence to answer this question. Key findings are as follows: (1) models are vulnerable to subtle visual noise and irrelevant distractors, failing to attend to salient objects; (2) when diagram–text conflicts arise, models over-rely on textual information, particularly those with weaker perceptual ability; and (3) models often resort to pattern memorization instead of perceptual understanding, as evidenced by their insensitivity to vertex ordering in shape classification, even though vertex order defines shape identity under formal geometric rules.
2. Does stronger perceptual ability lead to better reasoning performance? To understand the potential reasons for weak diagram perception in MLLMs, we examine the limitations of existing diagram–caption training datasets, such as MAVIS [20] and AutoGeo [21], two of the largest in mathematical vision. They often contain ambiguous expressions and obscure structural properties of diagrams (Fig. 2(a)). As a means of investigation, we construct GEOMETRIC, a high-quality {geometry-image, text} pairs that encode shapes, attributes, and interrelationships as graphs with fine-grained bounding box locations. Training on GEOMETRIC yields strong improvements in perception tasks, achieving +79% on the grounding task (Fig. 1). Additional results for Qwen2.5-VL-7B/32B trained on GEOMETRIC are presented in §3.2.
Moreover, in the close-up comparison with math-specific MLLMs (Fig. 2(b), colored dot points), training with our proposed dataset substantially improves perceptual performance over AutoGeo and MAVIS, by 23.0% and 35.6%, respectively. In contrast, two alternative variants underperform the baseline (SVE-Math-DeepSeek-7B), with MAVIS diagram–caption pairs (despite being 5 $\times$ larger in scale than ours) yielding poorer results due to high uncertainty and the out-of-distribution nature of real-world geometric diagrams.